the number of legislators has increased approximately linearly from 1137 positions in 2004 to 1265 positions in 2012. Let n be the number of legislators at t years since 2000. Find an equation of a linear model to describe the data.
Answer by TimothyLamb(4379)

slope = m = dy/dx
m = (1265 - 1137)/(2012 - 2004)
m = 16
---
linear equation for the years 2004 to 2012:
n = 16t + 1137
---
n for the year 2000:
n = 16t + 1137
n = 16*(2000-2004) + 1137
n = 16*(-4) + 1137
n = 1073
---
answer:
linear equation for the years 2000 to 2012:
n = 16t + 1073
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php